# Large Language Models for Information Retrieval: A Survey

### v1

@@ 1 Introduction
Large language models (LLMs) have recently emerged as transformative forces across various research fields, such as natural language processing (NLP) [33, 34, 35], recommender systems [36, 37, 38, 39], and even molecule discovery [40]. These cutting-edge LLMs are primarily based on the Transformer architecture and undergo extensive pre-training on diverse textual sources, including web pages, research articles, books, and codes. As their scale continues to expand (including both model size and data volume), LLMs have demonstrated remarkable advances in their capabilities. On the one hand, LLMs have exhibited unprecedented proficiency in language understanding and generation, resulting in responses that are more human-like and better align with human intentions. On the other hand, the larger LLMs have shown impressive emergent abilities when dealing with complex tasks [41], such as generalization and reasoning skills. Notably, LLMs can effectively apply their learned knowledge and reasoning abilities to tackle new tasks with just a few task-specific demonstrations or appropriate instructions [instructgpt, flan]. Furthermore, advanced techniques, such as in-context learning, have significantly enhanced the generalization performance of LLMs without requiring fine-tuning on specific downstream tasks [34]. This breakthrough is particularly valuable, as it reduces the need for extensive fine-tuning while attaining remarkable task performance. Powered by prompting strategies such as chain-of-thought, LLMs can generate outputs with step-by-step reasoning, navigating complex decision-making processes [42]. Leveraging the impressive power of LLMs can undoubtedly improve the performance of IR systems. By incorporating these sophisticated language models, IR systems can provide users with more accurate responses, ultimately reshaping the landscape of information access and retrieval.
@@ 2.1 Information Retrieval
The field of IR has witnessed significant advancement with the emergence of various models over time. One such early model is the Boolean model, which employs Boolean logic operators to combine query terms and retrieve documents that satisfy specific conditions [25]. Based on the “bag-of-words” assumption, the vector space model [26] represents documents and queries as sparse term-based vectors. Relevance estimation is then performed by assessing the lexical similarity between the sparse query and document vectors. The efficiency of this model is further improved through the effective organization of text content using the inverted index. Moving towards more sophisticated approaches, statistical language models have been introduced to estimate the likelihood of term occurrences and incorporate context information, leading to more accurate and context-aware retrieval [27, 51]. In recent years, the neural IR [30, 52, 53] paradigm has gained considerable attention in the research community. By harnessing the powerful representation capabilities of neural networks, this paradigm can capture semantic relationships between queries and documents, thereby significantly enhancing retrieval performance.
@@ 3.2.2 Inherent Knowledge of LLMs and Document Corpus
Although LLMs exhibit remarkable capabilities, their lack of familiarity with specific domains can lead to the generation of hallucinatory or irrelevant queries. To address this issue, recent studies [92, 93, 94, 95] have proposed a hybrid approach that enhancing LLM-base query rewriting methods with document corpus. For instance, LLMs may not possess immediate comprehension of the query “ADHD”, but the examination of retrieved documents can offer valuable insights indicating its association with a medical condition referred to as “Attention-Deficit/Hyperactivity Disorder”.
Why incorporating a document corpus? The integration of a document corpus offers several notable advantages. First, it provides domain-specific knowledge by fine-tuning the query generation process for specific subject areas, enabling a targeted and specialized approach to IR. Second, it ensures the grounding of queries in reliable and verifiable knowledge by incorporating factual information extracted from the corpus. Third, it incorporates up-to-date concepts by enriching queries with contemporary information, surpassing the knowledge contained within LLMs, and ensuring enhanced information richness and timeliness. Finally, it supports relevance by utilizing relevant documents as a supplementary resource to refine the query generation process, effectively reducing the generation of irrelevant information and improving the production of contextually relevant outputs.
@@
$\bullet$ Combining the query and documents in the input of LLMs. In the era of LLMs, incorporating instructions within the prompts is a flexible method for achieving specific functionalities. There are several studies that leverage this procedure. GRM [92] uses relevance-aware sample estimation for more accurate weighting of rewrite terms. For the query intent classification task, QUILL [93] demonstrates that retrieval augmentation of queries provides LLMs with valuable additional context, enabling improved understanding. It proposes a two-stage distill pipeline to transfer the knowledge in LLMs to small models in order to rewrite queries more efficiently. LameR [96] proposes a retrieve-rewrite-retrieve framework. It takes in BM25 as the retriever without using any annotated query-document pairs, which shows the great potential of LLMs in zero-shot retrieval. Besides, there are no strict rules about whether rewriters should appear before retrievers. InteR [94] proposes a framework supporting multi-turn interaction between search engines and LLMs, which allows search engines to expand knowledge in queries using LLM-generated knowledge collections and enables LLMs to enhance prompt formulation using documents retrieved by search engines.
$\bullet$ Combining generative relevance feedback (GRF) from LLMs with pseudo relevance feedback (PRF) from the document corpus. GRF provides external context not present in first-pass retrieval, whereas PRF restricts the query to the information contained within the target corpus. Consequently, combining generative and pseudo-relevance feedback ranking signals is an effective way to achieve the benefits of both feedback classes. Based on this motivation, a weighted fusion method (GRF+PRF) that combines the ranking signals of GRF and PRF is adopted to improve recall-oriented effectiveness [95].
@@ 3.3.1 Prompting Methods
$\bullet$ Few-shot prompting. Few-shot prompting, also known as in-context learning, involves providing the model with a limited set of examples or demonstrations related to the desired task or domain [97, 98, 90, 87]. These examples serve as a form of explicit instruction, allowing the model to adapt its language generation to the specific task or domain at hand. Query2Doc [87] prompts the model to write a passage that answers the query given some query-document pairs provided by the ranking dataset, such as MSMARCO [99] and NQ [100]. This work experiment on a single prompt. To further study the impact of different prompt designing,  [90] explores eight different prompts, such as promting LLMs to generate query expansion terms instead of entire pseudo documents, CoT prompting and so on. There are some illustrative prompts in Table I. This work conducts more experiments than Query2Doc, but experiments show that this prompt is not as effective as that in Query2Doc.
$\bullet$ Chain-of-thought prompting. CoT prompting [42] is a strategy that involves iterative prompting, where the model is provided with a sequence of instructions or partial outputs [97, 92, 90]. In conversational search, the process of query re-writing is multi-turn, which means query should be refined step-by-step with the interaction between search engine and users. This process is naturally coincided with CoT process. As shown in 4, engineers conduct the CoT process through adding some instructions during each turn, such as “Based on all previous turns, xxx”. While in ad-hoc search, there is only one-round in query re-writing. So CoT could only be accomplished in a simple and coarse way. For example, as shown in Table I, researchers add “Give the rationale before answering” in the instructions to prompt LLMs think deeply [90].
@@ 3.3.3 Knowledge Distillation Methods
Although LLM-based methods have demonstrated significant improvements in query rewriting tasks, their practical implementation for online deployment is hindered by the substantial latency caused by the computational requirements of LLMs. To mitigate this challenge, knowledge distillation has emerged as a prominent technique in the industry. In the QUILL [93] framework, a two-stage distillation method is proposed. This approach entails utilizing a retrieval-augmented LLM as the professor model, a vanilla LLM as the teacher model, and a lightweight BERT model as the student model. The professor model is trained on two extensive datasets, namely Orcas-I [103] and EComm [93], which are specifically curated for query intent understanding. Subsequently, a two-stage distillation process is employed to transfer knowledge from the professor model to the teacher model, followed by knowledge transfer from the teacher model to the student model. Empirical findings demonstrate that this knowledge distillation methodology surpasses the simple scaling up of model size from base to XXL, resulting in even more substantial improvements. In a “rewrite-retrieve-read” framework [91]. They first prompt an LLM to rewrite the queries, and then conduct retrieval-augmented reading. To enhance the effectiveness of the framework, a trainable rewriter, implemented as a small language model, is employed to further modify the search queries to align with the requirements of the frozen retriever and the LLM reader. To fine-tune the rewriter, a two-step training process is utilized. A supervised warm-up training is first conducted using pseudo data. The retrieve-then-read pipeline is then modeled as a reinforcement learning context, and the rewriter is trained as a policy model to maximize the reward obtained from the pipeline’s performance.
@@ 4.2 Employing LLMs to Enhance Model Architecture
Leveraging the excellent text encoding and decoding capabilities of LLMs, it is feasible to understand queries and documents with greater precision compared to earlier smaller-sized models [56]. Researchers have endeavored to utilize LLMs as the foundation for constructing advanced retrieval models. These methods can be broadly grouped into two categories, i.e., encoder-based retrievers and generative retrievers.
@@ 4.2.1 Encoder-based Retriever
In addition to the quantity and quality of the data, the representative capability of models also greatly influences the efficacy of retrievers. Inspired by the LLM’s excellent ability to model and comprehend natural language, some researchers [113, 114, 115] leverage LLMs as retrieval encoders and investigate the impact of model scale on retriever performance.
General Retriever. Since the effectiveness of retrievers primarily relies on the capability of text embedding, the evolution of text embedding models often has a significant impact on the progress of retriever development. In the era of LLMs, a pioneer work is made by OpenAI [113]. They view the neighboring pieces of text as positive pairs to facilitate the unsupervised pre-training of a series of text embedding models, denoted as cpt-text, with model sizes varying from 300M to 175B parameters. The experiments on the MS MARCO [99] and BEIR [112] datasets indicate that larger model scales have the potential to yield improved performance in unsupervised learning and transfer learning for text search tasks. Nevertheless, pre-training LLMs from scratch is prohibitively expensive for most researchers. To overcome this limitation, T5-family models with smaller sizes, such as Base, Large, XL, and XXL, are used to initialize the model parameters of bi-encoders, which are then fine-tuned on retrieval datasets [114]. The experimental results confirm again that larger model sizes can lead to better performance, particularly in zero-shot settings.
@@ 5 Reranker
Reranker, as the second-pass document filter in IR, aims to rerank a set of documents retrieved by the retriever (e.g., BM25) based on their relevance to the user query. With respect to the application of LLMs, the existing reranking methods can be divided into three primary paradigms: fine-tuning LLMs for reranking, prompting LLMs for reranking, and utilizing LLMs for training data augmentation. These paradigms will be elaborated upon in the following sections. Recall that we will use the term document to refer to the text retrieved in general IR scenarios, including instances such as passages (e.g., passages in MS MARCO passage ranking dataset [99]).
@@ 5.1.1 Fine-tuning LLMs as Generation Models
In this field, existing studies usually formulate document ranking as a generation task and optimize LLMs by generative loss [13, 121, 122]. Specifically, reranking models are usually fine-tuned to generate a single token, such as “true” or “false”, given the query and documents. During inference, the query-document relevance score is determined based on the logit value of the generated token. For example, a T5 model can be fine-tuned to generate classification tokens for relevant or irrelevant query-document pairs [13]. At inference time, a softmax function is applied to the logits of “true” and “false” tokens, and the relevance score is computed as the probability assigned to the “true” token. The following method [121] involves a multi-view learning approach based on the T5 model. This approach simultaneously considers two tasks: generating classification tokens for a given query-document pair and generating the corresponding query conditioned on the provided document. DuoT5 [122] considers a triple $(q,d_{i},d_{j})$ as the input of the T5 model and is fine-tuned to generate token “true” if document $d_{i}$ is more relevant to query $q_{i}$ than document $d_{j}$, and “false” otherwise. During inference, for each document $d_{i}$, it enumerates all other documents $d_{j}$ and uses global aggregation functions to generate the relevance score $s_{i}$ for document $d_{i}$ (e.g., $s_{i}=\sum_{j}p_{i,j}$, where $p_{i,j}$ represents the probability of generating “true” when taking $(q,d_{i},d_{j})$ as the model input).
@@ 5.2.1 Pointwise methods
The pointwise methods measure the relevance between a single document and a query, and can be categorized into two types: relevance generation [125] and query generation [126, 127]. The upper part in Figure 6 (a) shows the relevance generation based on a given prompt, where LLMs output “Yes” if the document is relevant to the query and “No” otherwise. The probability that LLMs generate the token “Yes” or “No” is used to calculate the query-document relevance score $s$ as follows:
@@
where $|q|$ denotes the token number of query $q$, $d$ denotes the document, and $\mathcal{P}$ represents the provided prompt. The documents are then ranked according to their relevance scores. Existing query generation-based methods [126] primarily rely on a handcrafted prompt (e.g., “Please write a query based on this document”), which may not be optimal. As prompt is a key factor in instructing LLMs to perform various NLP tasks, it is important to identify the optimal prompt for best model performance. Along this line, a discrete prompt optimization method Co-Prompt [127] is proposed for better prompt generation in reranking tasks.

### v2

@@ 1 Introduction
Large language models (LLMs) have recently emerged as transformative forces across various research fields, such as natural language processing (NLP) [33, 34, 35], recommender systems [36, 37, 38, 39], finance [40], and even molecule discovery [41]. These cutting-edge LLMs are primarily based on the Transformer architecture and undergo extensive pre-training on diverse textual sources, including web pages, research articles, books, and codes. As their scale continues to expand (including both model size and data volume), LLMs have demonstrated remarkable advances in their capabilities. On the one hand, LLMs have exhibited unprecedented proficiency in language understanding and generation, resulting in responses that are more human-like and better align with human intentions. On the other hand, the larger LLMs have shown impressive emergent abilities when dealing with complex tasks [42], such as generalization and reasoning skills. Notably, LLMs can effectively apply their learned knowledge and reasoning abilities to tackle new tasks with just a few task-specific demonstrations or appropriate instructions [43, 44]. Furthermore, advanced techniques, such as in-context learning, have significantly enhanced the generalization performance of LLMs without requiring fine-tuning on specific downstream tasks [34]. This breakthrough is particularly valuable, as it reduces the need for extensive fine-tuning while attaining remarkable task performance. Powered by prompting strategies such as chain-of-thought, LLMs can generate outputs with step-by-step reasoning, navigating complex decision-making processes [45]. Leveraging the impressive power of LLMs can undoubtedly improve the performance of IR systems. By incorporating these sophisticated language models, IR systems can provide users with more accurate responses, ultimately reshaping the landscape of information access and retrieval.
@@ 2.1 Information Retrieval
The field of IR has witnessed significant advancement with the emergence of various models over time. One such early model is the Boolean model, which employs Boolean logic operators to combine query terms and retrieve documents that satisfy specific conditions [25]. Based on the “bag-of-words” assumption, the vector space model [26] represents documents and queries as vectors in term-based space. Relevance estimation is then performed by assessing the lexical similarity between the query and document vectors. The efficiency of this model is further improved through the effective organization of text content using the inverted index. Moving towards more sophisticated approaches, statistical language models have been introduced to estimate the likelihood of term occurrences and incorporate context information, leading to more accurate and context-aware retrieval [27, 54]. In recent years, the neural IR [30, 55, 56] paradigm has gained considerable attention in the research community. By harnessing the powerful representation capabilities of neural networks, this paradigm can capture semantic relationships between queries and documents, thereby significantly enhancing retrieval performance.
@@ 3.2.2 Inherent Knowledge of LLMs and Document Corpus
Although LLMs exhibit remarkable capabilities, their lack of familiarity with specific domains can lead to the generation of hallucinatory or irrelevant queries. To address this issue, recent studies [95, 96, 97, 98] have proposed a hybrid approach that enhancing LLM-base query rewriting methods with document corpus.
Why incorporating a document corpus? The integration of a document corpus offers several notable advantages. First, it provides domain-specific knowledge by fine-tuning the query generation process for specific subject areas, enabling a targeted and specialized approach to IR. For instance, LLMs may not possess immediate comprehension of the query “ADHD”, but the examination of retrieved documents can offer valuable insights indicating its association with a medical condition referred to as “Attention-Deficit/Hyperactivity Disorder”. Second, it ensures the grounding of queries in reliable and verifiable knowledge by incorporating factual information extracted from the corpus. Third, it incorporates up-to-date concepts by enriching queries with contemporary information, surpassing the knowledge contained within LLMs, and ensuring enhanced information richness and timeliness. Finally, it supports relevance by utilizing relevant documents as a supplementary resource to refine the query generation process, effectively reducing the generation of irrelevant information and improving the production of contextually relevant outputs.
@@
$\bullet$ Late fusion of LLM-only re-writing and traditional PRF retrieval results. Traditional PRF methods leverage relevant documents retrieved from a document corpus to rewrite queries, which restricts the query to the information contained in the target corpus. On the contrary, LLM-only re-writing methods provides external context not present in the corpus, which is more diverse. Both approaches have potential to independently enhance retrieval performance. Therefore, a straightforward strategy for combining them is using a weighted fusion method on retrieval results, which is demonstrated effective [98].
$\bullet$ Combining retrieved relevant documents in the prompts of LLMs. In the era of LLMs, incorporating instructions within the prompts is a flexible method for achieving specific functionalities. There are several studies that leverage this procedure. For the query intent classification task, QUILL [96] demonstrates that retrieval augmentation of queries provides LLMs with meaningful supportive context, which leads to better understanding. It proposes a two-stage distill pipeline to transfer the knowledge in LLMs to small models in order to rewrite queries more efficiently. LameR [99] proposes a retrieve-rewrite-retrieve framework. It takes in BM25 as the retriever without using any annotated query-document pairs, which shows the great potential of LLMs in zero-shot retrieval. Besides, there are no strict rules about whether rewriters should appear before retrievers. InteR [97] proposes a framework supporting multi-turn interaction between search engines and LLMs. This setup enables search engines to expand queries with LLM-generated knowledge, while simultaneously permitting LLMs to refine prompt formulations with relevant documents provided by search engines.
$\bullet$ Enhancing factuality of generative relevance feedback (GRF) by pseudo relevance feedback (PRF). Although generative documents are often relevant and diverse, they exhibit hallucinatory characteristics. In contrast, traditional documents are generally regarded as reliable sources of factual information. Motivated by this observation, GRM [95] proposes a novel technique known as relevance-aware sample estimation (RASE). RASE leverages relevant documents retrieved from the collection to assign weights to generated documents. In this way, GRM ensures that relevance feedback is not only diverse but also maintain a high degree of factuality.
@@ 3.3.1 Prompting Methods
$\bullet$ Few-shot prompting. Few-shot prompting, also known as in-context learning, involves providing the model with a limited set of examples or demonstrations related to the desired task or domain [100, 101, 93, 86]. These examples serve as a form of explicit instruction, allowing the model to adapt its language generation to the specific task or domain at hand. Query2Doc [86] prompts LLMs to write a document that answers the query with some demo query-document pairs provided by the ranking dataset, such as MSMARCO [102] and NQ [103]. This work experiment on a single prompt. To further study the impact of different prompt designing,  [93] explores eight different prompts, such as promting LLMs to generate query expansion terms instead of entire pseudo documents, CoT prompting and so on. There are some illustrative prompts in Table I. This work conducts more experiments than Query2Doc, but experiments show that this prompt is not as effective as that in Query2Doc.
$\bullet$ Chain-of-thought prompting. CoT prompting [45] is a strategy that involves iterative prompting, where the model is provided with a sequence of instructions or partial outputs [100, 93]. In conversational search, the process of query re-writing is multi-turn, which means query should be refined step-by-step with the interaction between search engine and users. This process is naturally coincided with CoT process. As shown in 4, engineers conduct the CoT process through adding some instructions during each turn, such as “Based on all previous turns, xxx”. While in ad-hoc search, there is only one-round in query re-writing. So CoT could only be accomplished in a simple and coarse way. For example, as shown in Table I, researchers add “Give the rationale before answering” in the instructions to prompt LLMs think deeply [93].
@@ 3.3.3 Knowledge Distillation Methods
Although LLM-based methods have demonstrated significant improvements in query rewriting tasks, their practical implementation for online deployment is hindered by the substantial latency caused by the computational requirements of LLMs. To mitigate this challenge, knowledge distillation has emerged as a prominent technique in the industry. In the QUILL [96] framework, a two-stage distillation method is proposed. This approach entails utilizing a retrieval-augmented LLM as the professor model, a vanilla LLM as the teacher model, and a lightweight BERT model as the student model. The professor model is trained on two extensive datasets, namely Orcas-I [106] and EComm [96], which are specifically curated for query intent understanding. Subsequently, a two-stage distillation process is employed to transfer knowledge from the professor model to the teacher model, followed by knowledge transfer from the teacher model to the student model. Empirical findings demonstrate that this knowledge distillation methodology surpasses the simple scaling up of model size from base to XXL, resulting in even more substantial improvements. In a recently proposed “rewrite-retrieve-read” framework [94], an LLM is first used to rewrite the queries by prompting, followed by a retrieval-augmented reading process. To improve framework effectiveness, a trainable rewriter, implemented as a small language model, is incorporated to further adapt search queries to align with both the frozen retriever and the LLM reader’s requirements. The rewriter’s refinement involves a two-step training process. Initially, supervised warm-up training is conducted using pseudo data. Subsequently, the retrieve-then-read pipeline is described as a reinforcement learning scenario, with the rewriter’s training acting as a policy model to maximize pipeline performance rewards.
@@ 4.2 Employing LLMs to Enhance Model Architecture
Leveraging the excellent text encoding and decoding capabilities of LLMs, it is feasible to understand queries and documents with greater precision compared to earlier smaller-sized models [59]. Researchers have endeavored to utilize LLMs as the foundation for constructing advanced retrieval models. These methods can be grouped into two categories, i.e., encoder-based retrievers and generative retrievers.
@@ 4.2.1 Encoder-based Retriever
In addition to the quantity and quality of the data, the representative capability of models also greatly influences the efficacy of retrievers. Inspired by the LLM’s excellent capability to encode and comprehend natural language, some researchers [115, 116, 117] leverage LLMs as retrieval encoders and investigate the impact of model scale on retriever performance.
General Retriever. Since the effectiveness of retrievers primarily relies on the capability of text embedding, the evolution of text embedding models often has a significant impact on the progress of retriever development. In the era of LLMs, a pioneer work is made by OpenAI [115]. They view the adjacent text segments as positive pairs to facilitate the unsupervised pre-training of a set of text embedding models, denoted as cpt-text, whose parameter values vary from 300M to 175B. Experiments conducted on the MS MARCO [102] and BEIR [114] datasets indicate that larger model scales have the potential to yield improved performance in unsupervised learning and transfer learning for text search tasks. Nevertheless, pre-training LLMs from scratch is prohibitively expensive for most researchers. To overcome this limitation, T5-family models with smaller sizes, such as Base, Large, XL, and XXL, are used to initialize the model parameters of bi-encoders, which are then fine-tuned on retrieval datasets [116]. The experimental results confirm again that larger model sizes can lead to better performance, particularly in zero-shot settings.
@@ 5 Reranker
Reranker, as the second-pass document filter in IR, aims to rerank a document list retrieved by the retriever (e.g., BM25) based on the query-document relevance. With respect to the application of LLMs, the existing reranking methods can be divided into three primary paradigms: fine-tuning LLMs for reranking, prompting LLMs for reranking, and utilizing LLMs for training data augmentation. These paradigms will be elaborated upon in the following sections. Recall that we will use the term document to refer to the text retrieved in general IR scenarios, including instances such as passages (e.g., passages in MS MARCO passage ranking dataset [102]).
@@ 5.1.1 Fine-tuning LLMs as Generation Models
In this field, existing studies usually formulate document ranking as a generation task and optimize LLMs by generative loss [13, 123, 124]. Specifically, reranking models are usually fine-tuned to generate a single token, such as “true” or “false”, given the query and documents. During inference, the query-document relevance score is determined based on the logit of the generated token. For example, a T5 model can be fine-tuned to generate classification tokens for relevant or irrelevant query-document pairs [13]. At inference time, a softmax function is applied to the logits of “true” and “false” tokens, and the relevance score is calculated as the probability of the “true” token. The following method [123] involves a multi-view learning approach based on the T5 model. This approach simultaneously considers two tasks: generating classification tokens for a given query-document pair and generating the corresponding query conditioned on the provided document. DuoT5 [124] considers a triple $(q,d_{i},d_{j})$ as the input of the T5 model and is fine-tuned to generate token “true” if document $d_{i}$ is more relevant to query $q_{i}$ than document $d_{j}$, and “false” otherwise. During inference, for each document $d_{i}$, it enumerates all other documents $d_{j}$ and uses global aggregation functions to generate the relevance score $s_{i}$ for document $d_{i}$ (e.g., $s_{i}=\sum_{j}p_{i,j}$, where $p_{i,j}$ represents the probability of generating “true” when taking $(q,d_{i},d_{j})$ as the model input).
@@ 5.2.1 Pointwise methods
The pointwise methods measure the relevance between a query and a single document, and can be categorized into two types: relevance generation [127] and query generation [128, 129]. The upper part in Figure 6 (a) shows the relevance generation based on a given prompt, where LLMs output “Yes” if the document is relevant to the query and “No” otherwise. The probability that LLMs generate the token “Yes” or “No” is used to calculate the query-document relevance score $s$ as follows:
@@
where $|q|$ denotes the token number of query $q$, $d$ denotes the document, and $\mathcal{P}$ represents the provided prompt. The documents are then reranked based on their relevance scores. Existing query generation-based methods [128] primarily rely on a handcrafted prompt (e.g., “Please write a query based on this document”), which may not be optimal. As prompt is a key factor in instructing LLMs to perform various NLP tasks, it is important to identify the optimal prompt for best model performance. Along this line, a discrete prompt optimization method Co-Prompt [129] is proposed for better prompt generation in reranking tasks.